# Spin transport and lack of quantisation in the $A\mathrm{II}$ class on the honeycomb structure

### v1

@@ 3.2. Linear response à la Kubo
We are interested in studying the linear response coefficients of insulators in the $\mathrm{AII}^{\hexagon}$ class, subject to an external homogeneous electric field which is small in its intensity and switched on adiabatically in time. In previous work [MaPaTe], this goal was accomplished via the construction of the so-called non-equilibrium almost-stationary state (NEASS). Although here we borrow some ideas from [MaPaTe], by exploiting the symmetries of the class $\mathrm{AII}^{\hexagon}$ we are able to employ the standard Kubo approach [Ku], avoiding the construction of the NEASS.
@@
The following lemma, whose proof can be found in [AG, AW], will be repeatedly used.
@@
Let $(H,\mu)$ be an insulator. If $H\in\mathcal{P}_{0}(\mathcal{H})$, then $P\in\mathcal{P}_{0}(\mathcal{H})$.
@@
where $\mathcal{D}(X_{i})$ denotes its maximal dense domain. We will consider commutators of bounded operators with the position operator, for which the following standard result holds (here we recall it for completeness).
@@
Let $A$ be in $\mathcal{B}_{0}(\mathcal{H})$ (resp. in $\mathcal{P}_{0}(\mathcal{H})$). Then, $[A,X_{i}]$ is in $\mathcal{B}_{0}(\mathcal{H})$ (resp. in $\mathcal{P}_{0}(\mathcal{H})$).
@@
Recalling the dimerization notation in (2.4), we obtain that
@@
thus $[A,X_{i}]\in\mathcal{B}_{0}(\mathcal{H})$. Finally, if $A$ is periodic, by the Jacobi identity for commutators we get
@@
where we have used that $[X_{i},T_{\gamma}]=\gamma_{i}T_{\gamma}$. ∎
@@
which shows that, even though $[X_{j},P]$ is bounded (see Lemma 3.5), computing the adiabatic limit $\eta\to 0^{+}$ of the term $L_{\eta,j}$ is non-trivial. As it turns out, the fact that $(H,\mu)$ is an insulator has to be used in an essential way in order to perform such a limit. In preparation for the next results, it is convenient to introduce the following notation: for any linear operator $A$ on $\mathcal{H}$ (we will consider operators in $\mathcal{B}(\mathcal{H})$ or the position operator $X$), such that $AP$ is densely defined, we denote its diagonal and off-diagonal parts, with respect to the orthogonal projection $P$, by
@@
Observe that $[X_{j},P]\in\mathcal{P}_{0}(\mathcal{H})$ by Lemma 3.5 and Lemma 3.4. Note that the operator $[X_{j},P]$ is off-diagonal with respect to the decomposition induced by the projection $P$, see (3.10). Therefore, Lemma 3.7(i) implies that $I([X_{j},P])\equiv P_{1}$ exists in $\mathcal{P}(\mathcal{H}).$
@@
    Let us observe that the operator $P_{1}$ coincides with $\Pi_{1}$ in [MaPaTe, Proposition 4.1.2]. Indeed, both of these operators are defined by computing the inverse of the Liouvillian $I$ on $[X_{j},P]$, with the difference that in the current paper the super-operator $I$ is written in terms of a time integral over the real line, while in [MaPaTe] the map $I$ is given by an energy integral in the complex plane (see e. g. [MaPaTe, Equation (4.2)]). Thus, in particular we have the following identity
@@ 3.3. Spin conductivity
$J_{i}^{\mathrm{conv}}$ $:=\frac{\mathrm{i}}{2}[H,X_{i}]S_{z}+\frac{\mathrm{i}}{2}S_{z}[H,X_{i}]$ (3.16)
$J_{i}^{\mathrm{prop}}$ $:=\mathrm{i}[H,X_{i}\,S_{z}]=X_{i}\,\mathscr{T}_{z}+\mathrm{i}[H,X_{i}]\,S_{z},$
@@
Note that if the spin $S_{z}$ is a conserved quantity, that is, $[H,S_{z}]=0$, the two definitions basically coincide, $J_{i}^{\mathrm{conv}}\equiv J_{i}^{\mathrm{prop}}$. However, when spin is not a conserved quantity, associating a current to it becomes ambiguous, and this remains a topic of debate in condensed matter physics. The first definition has been adopted by several works, such as [SCNSJM, SSTH, Sch], while the second, more recent definition was proposed by [SZXN]. Both definitions come with their own advantages and disadvantages, depending also on the context in which they are applied. For the moment, we point out that $J_{i}^{\mathrm{prop}}$ presents more technical challenges: whenever $H\in\mathcal{P}_{0}(\mathcal{H})$, then $J_{i}^{\mathrm{conv}}\in\mathcal{P}_{0}(\mathcal{H})$ by Lemma 3.5, while $J_{i}^{\mathrm{prop}}$ is neither periodic nor bounded because of the presence of the term $\mathrm{i}X_{i}\,\mathscr{T}_{z}$. In our analysis we follow [MaPaTe] and consider both definitions of the spin current.
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}(\mathcal{H})\cap\mathrm{AII}^{\hexagon}$. The conventional and proper spin conductivities, respectively denoted by $\sigma^{\mathrm{s},\mathrm{conv}}_{ij}$ and $\sigma^{\mathrm{s},\mathrm{prop}}_{ij}$, are defined as
@@
While $\sigma^{\mathrm{s},\mathrm{conv}}_{ij}$ is clearly well-posed, see Remark 3.10(i), this is not obvious for $\sigma^{\mathrm{s},\mathrm{prop}}_{ij}$, given that $J_{i}^{\mathrm{prop}}$ is an unbounded and non-periodic operator. Well-posedness of the latter is shown in the following theorem, which is the main result of this section.
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}(\mathcal{H})\cap\mathrm{AII}^{\hexagon}$ and let $P=P(H,\mu)$ be its Fermi projection. Let $\sigma^{\mathrm{s},\sharp}_{ij}=\sigma^{\mathrm{s},\sharp}_{ij}(H,\mu)$ be the spin conductivity as defined in (3.18), with $\sharp\in\{\mathrm{conv},\mathrm{prop}\}$. Then, we have that:
@@
    In particular, in the spin-commuting case, i. e. $[H,S_{z}]=0$, we have that
@@
Thus, for every insulator $(H,\mu)$ with $H\in\mathcal{P}_{0}(\mathcal{H})\cap\mathrm{AII}^{\hexagon}$ we are allowed to adopt the notation
@@
Before proving Proposition 3.13, we establish an important geometrical property of the spin conductivity, namely that it is an antisymmetric tensor. This property makes the spin conductivity invariant under rotations, and thus establishes that it measures an intrinsic transverse response, the spin Hall response, regardless of the orientation of the laboratory [SVWBJ]. Unlike the case of charge conductivity, this property is not obvious from its definition.
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}(\mathcal{H})\cap\mathrm{AII}^{\hexagon}$. The spin conductivity $\sigma^{\mathrm{s}}_{ij}(H,\mu)$ is an antisymmetric tensor, that is
@@
Since $\sigma^{\mathrm{s}}_{ij}=\lim_{\eta\to 0^{+}}\tau(J_{i}^{\mathrm{conv}}L_{\eta,j})$, it suffices to prove that the quantity $\tau(J_{i}^{\mathrm{conv}}L_{\eta,j})$ is antisymmetric. First of all, we show $\tau(J_{i}^{\mathrm{conv}}L_{\eta,i})=0$. This is a simple consequence of the symmetry under $v\otimes h$ reflection, see (2.13). Indeed, since
$R_{v\otimes h}HR_{v\otimes h}^{-1}=H,\qquad R_{v\otimes h}X_{i}R_{v\otimes h}^{-1}=(-)^{i-1}X_{i},\qquad R_{v\otimes h}S_{z}R_{v\otimes h}^{-1}=-S_{z}$
we have
$R_{v\otimes h}J_{i}^{\mathrm{conv}}R_{v\otimes h}^{-1}=(-)^{i}J_{i}^{\mathrm{conv}}\;,\qquad R_{v\otimes h}L_{\eta,j}R_{v\otimes h}^{-1}=(-)^{j-1}L_{\eta,j}\;.$
Therefore, since $J_{i}^{\mathrm{conv}}L_{\eta,j}\in\mathcal{P}(\mathcal{H})$ by using Lemma 3.2(iv) and the identity $R_{v\otimes h}\chi_{1}R^{-1}_{v\otimes h}=\chi_{1}$, we conclude that
$\begin{split}\tau(J_{i}^{\mathrm{conv}}L_{\eta,i})=\tau(R_{v\otimes h}J_{i}^{\mathrm{conv}}L_{\eta,i}R_{v\otimes h}^{-1})=-\tau(J_{i}^{\mathrm{conv}}L_{\eta,i})\;.\end{split}$
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}(\mathcal{H})\cap\mathrm{AII}^{\hexagon}$ and let $P=P(H,\mu)$ be its Fermi projector. Let $\mathscr{T}_{z}$ and $L_{\eta,j}$ be the operators defined respectively in (3.17) and (3.9). Then, we have that:
@@
(i) & (ii) Note that the operator $J_{i}^{\mathrm{conv}}\,L_{\eta,j}$ is periodic and bounded since both $J_{i}^{\mathrm{conv}}$ and $[P,X_{j}]$ are short-range by Lemmas 3.4 and 3.5. Thus, Lemma 3.2(i) ensures that $\sigma_{ij}^{\mathrm{s},\mathrm{conv}}$ is well-defined. On the other hand, for the proper spin conductivity $\sigma_{ij}^{\mathrm{s},\mathrm{prop}}$ we have that $\tau(X_{i}\mathscr{T}_{z}\,L_{\eta,j})=0$ in view of Proposition 3.16(ii), thus only $\mathrm{i}[H,X_{i}]S_{z}L_{\eta,j}\in\mathcal{P}(\mathcal{H})$ contributes to its value. Specifically, we get that
@@
To this end, it is convenient to rewrite the argument of the real part in (3.22) by using that the operator $L_{\eta,j}$ is off-diagonal due to the off-diagonality of $[X_{j},P]$. By exploiting the cyclicity of the trace per unit volume, we have that
@@
where we have used repeatedly the cyclicity of the trace per unit volume, and in the last equality we have exploited that $[S_{z},[X_{j},P]]=-[P,[S_{z},X_{j}]]-[X_{j},[P,S_{z}]]=[[P,S_{z}],X_{j}]$. By taking the real part of the second term on the last line of (3.24), we obtain the term in (3.19); while in the last line of (3.24) the real part of the first summand gives the third term in the expression for $\Sigma^{\mathrm{snc}}_{ij}$, see (3.20). Next, we collect the remaining terms contributing to the spin conductivity:
@@
(iii) follows immediately from (ii) by using that $\Sigma^{\mathrm{snc}}_{ij}=0$ due to $[H,S_{z}]=0=[P,S_{z}]$. ∎
@@
Since Remark 3.10(ii) ensures that the linear response operator $P_{1}$ computed via Kubo’s formula agrees with $\Pi_{1}$ given by the NEASS approach in [MaPaTe], it is clear that a priori these two methods yield the same formulas for the spin conductivies. Here, we have preferred to rewrite the spin-commuting term $\Sigma^{\mathrm{sc}}_{ij}$ in (3.19) by having the operator $S_{z}$ at the first place in the argument of the trace per unit volume, while in [MaPaTe, Theorem 5.6] $S_{z}$ appears nested in a commutator involving $[X_{i},P]$ and $[X_{j},P]$.

### v2

@@ 3.2. Linear response à la Kubo
We are interested in studying the linear response coefficients of insulators in $\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$, subject to an external homogeneous electric field which is small in its intensity and switched on adiabatically in time. In previous work [MaPaTe], this goal was accomplished via the construction of the so-called non-equilibrium almost-stationary state (NEASS). Although here we borrow some ideas from [MaPaTe], by exploiting the spatial symmetries we are able to employ the standard Kubo approach [Ku], avoiding the construction of the NEASS.
@@
The following lemma, whose proof can be found in [AG, AW], see also Lemma 3.5(i) below, will be repeatedly used.
@@
Let $(H,\mu)$ be an insulator. If $H\in\mathcal{P}_{0}(\mathcal{H})$, then $P\in\mathcal{P}_{0}(\mathcal{H})$ and $(\mathbb{1}+\mathrm{e}^{\beta(H-\mu)})^{-1}\in\mathcal{P}_{0}(\mathcal{H})$ for any $\beta>0$.
@@
where $\mathcal{D}(X_{i})$ denotes its maximal dense domain. The following standard result holds.
@@
  1. (i)
    Let $A,B\in\mathcal{P}_{0}(\mathcal{H})$ then $AB\in\mathcal{P}_{0}(\mathcal{H})$. Furthermore, if $f$ is analytic in a strip around the real axis, then $f(A)\in\mathcal{P}_{0}(\mathcal{H})$.
  2. (ii)
    Let $A$ be in $\mathcal{B}_{0}(\mathcal{H})$ (resp. in $\mathcal{P}_{0}(\mathcal{H})$). Then, $A,X_{i}]$ is in $\mathcal{B}_{0}(\mathcal{H})$ (resp. in $\mathcal{P}_{0}(\mathcal{H})$).
@@
The first claim in (i) follows by the triangle inequality. To prove the second claim in (i), we note that by Cauchy integral formula we can write
$f(H)=\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\mathrm{d}zf(z)(H-z)^{-1}\;,$
where $\mathcal{C}$ is a positively-oriented complex contour within the strip of analyticity of $f$ and crossing the real axis away from the spectrum of $H$. The claim then follows by Combes–Thomas estimates for the resolvent $(H-z)^{-1}$, which give exponential decay for such $z\in\mathcal{C}$.
We now prove (ii). Recalling the dimerization notation in (2.4), we obtain that
@@
thus $A,X_{i}]\in\mathcal{B}_{0}(\mathcal{H})$. Finally, if $A$ is periodic, by the Jacobi identity for commutators we get
@@
where we have used that $X_{i},T_{\gamma}]=\gamma_{i}T_{\gamma}$. ∎
@@
which shows that, even though $X_{j},P]$ is bounded (see Lemma 3.5), computing the adiabatic limit $\eta\to 0^{+}$ of the term $L_{\eta,j}$ is non-trivial. As it turns out, the fact that $(H,\mu)$ is an insulator has to be used in an essential way in order to perform such a limit. In preparation for the next results, it is convenient to introduce the following notation: for any linear operator $A$ on $\mathcal{H}$ (we will consider operators in $\mathcal{B}(\mathcal{H})$ or the position operator $X$), such that $AP$ is densely defined, we denote its diagonal and off-diagonal parts, with respect to the orthogonal projection $P$, by
@@
Observe that $X_{j},P]\in\mathcal{P}_{0}(\mathcal{H})$ by Lemma 3.5 and Lemma 3.4. Note that the operator $X_{j},P]$ is off-diagonal with respect to the decomposition induced by the projection $P$, see (3.10). Therefore, Lemma 3.7(i) implies that $I([X_{j},P])\equiv P_{1}$ exists in $\mathcal{P}(\mathcal{H}).$
@@
    Let us observe that the operator $P_{1}$ coincides with $\Pi_{1}$ in [MaPaTe, Proposition 4.1.2]. Indeed, both of these operators are defined by computing the inverse of the Liouvillian $I$ on $X_{j},P]$, with the difference that in the current paper the super-operator $I$ is written in terms of a time integral over the real line, while in [MaPaTe] the map $I$ is given by an energy integral in the complex plane (see e. g. [MaPaTe, Equation (4.2)]). Thus, in particular we have the following identity
@@ 3.3. Spin conductivity
$J_{i}^{\mathrm{conv}}$ $=\frac{\mathrm{i}}{2}[H,X_{i}]S_{z}+\frac{\mathrm{i}}{2}S_{z}[H,X_{i}]$ (3.16)
$J_{i}^{\mathrm{prop}}$ $=\mathrm{i}[H,X_{i}\,S_{z}]=X_{i}\,\mathscr{T}_{z}+\mathrm{i}[H,X_{i}]\,S_{z},$
@@
Note that if the spin $S_{z}$ is a conserved quantity, that is, $H,S_{z}]=0$, the two definitions basically coincide, $J_{i}^{\mathrm{conv}}\equiv J_{i}^{\mathrm{prop}}$. However, when spin is not a conserved quantity, associating a current to it becomes ambiguous, and this remains a topic of debate in condensed matter physics. The first definition has been adopted by several works, such as [SCNSJM, SSTH, Sch], while the second, more recent definition was proposed by [SZXN]. Both definitions come with their own advantages and disadvantages, depending also on the context in which they are applied. For the moment, we point out that $J_{i}^{\mathrm{prop}}$ presents more technical challenges: whenever $H\in\mathcal{P}_{0}(\mathcal{H})$, then $J_{i}^{\mathrm{conv}}\in\mathcal{P}_{0}(\mathcal{H})$ by Lemma 3.5, while $J_{i}^{\mathrm{prop}}$ is neither periodic nor bounded because of the presence of the term $X_{i}\,\mathscr{T}_{z}$. In our analysis we follow [MaPaTe] and consider both definitions of the spin current.
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$. The conventional and proper spin conductivities, respectively denoted by $\sigma^{\mathrm{s},\mathrm{conv}}_{ij}$ and $\sigma^{\mathrm{s},\mathrm{prop}}_{ij}$, are defined as
@@
While $\sigma^{\mathrm{s},\mathrm{conv}}_{ij}$ is clearly well-posed, see Remark 3.10(i), this is not obvious for $\sigma^{\mathrm{s},\mathrm{prop}}_{ij}$, given that $J_{i}^{\mathrm{prop}}$ is an unbounded and non-periodic operator. Well-posedness of the latter is shown in the following proposition, which is the main result of this section.
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$. Let $P=P(H,\mu)$ be its Fermi projection. Let $\sigma^{\mathrm{s},\sharp}_{ij}=\sigma^{\mathrm{s},\sharp}_{ij}(H,\mu)$ be the spin conductivity as defined in (3.18), with $\sharp\in\{\mathrm{conv},\mathrm{prop}\}$. Then, we have that:
@@
    In particular, in the spin-commuting case, i. e. $H,S_{z}]=0$, we have that
@@
Thus, for every insulator $(H,\mu)$ with $H\in\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$ we are allowed to adopt the notation
@@
Before proving Proposition 3.13, we establish an important geometric property of the spin conductivity, namely that it is an antisymmetric tensor under further symmetry assumptions. This property makes the spin conductivity invariant under rotations, and thus establishes that it measures an intrinsic transverse response, the spin Hall response, regardless of the orientation of the laboratory [SVWBJ]. Unlike the case of charge conductivity, this property is not obvious from its definition and follows from using both $2\pi/3$-rotation and a (spatial) inversion symmetry.
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$ inversion symmetric. The spin conductivity $\sigma^{\mathrm{s}}_{ij}(H,\mu)$ is an antisymmetric tensor, that is
@@
Since $\sigma^{\mathrm{s}}_{ij}=\lim_{\eta\to 0^{+}}\tau(J_{i}^{\mathrm{conv}}L_{\eta,j})$, it suffices to prove that the quantity $\tau(J_{i}^{\mathrm{conv}}L_{\eta,j})$ is antisymmetric. First of all, we show $\tau(J_{i}^{\mathrm{conv}}L_{\eta,i})=0$. This is a simple consequence of the symmetry under inversion. Indeed, since
$\Pi_{a}\,H\,\Pi_{a}^{-1}=H,\qquad\Pi_{a}\,X_{i}\,\Pi_{a}^{-1}=(-)^{i+a}X_{i},\qquad\Pi_{a}\,S_{z}\,\Pi_{a}^{-1}=-S_{z}$
we have (5) (5) (5)Since $J_{i}^{\mathrm{conv}}$ and $L_{\eta,j}$ involve commutators with the position operator, their transformation properties are independent of the choice of the origin; see discussion before (2.14).
$\Pi_{a}\,J_{i}^{\mathrm{conv}}\,\Pi_{a}^{-1}=(-)^{i+a+1}J_{i}^{\mathrm{conv}}\;,\qquad\Pi_{a}\,L_{\eta,j}\,\Pi_{a}^{-1}=(-)^{j+a}L_{\eta,j}\;.$
Therefore, since $J_{i}^{\mathrm{conv}}L_{\eta,j}\in\mathcal{P}(\mathcal{H})$ by using Lemma 3.2(iv) and the identity $\Pi_{a}\,\chi_{1}\,\Pi_{a}^{-1}=\chi_{1}$, we conclude that
$\begin{split}\tau(J_{i}^{\mathrm{conv}}L_{\eta,i})=\tau(\Pi_{a}\,J_{i}^{\mathrm{conv}}L_{\eta,i}\,\Pi_{a}^{-1})=-\tau(J_{i}^{\mathrm{conv}}L_{\eta,i})\;.\end{split}$
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$. Let $P=P(H,\mu)$ be its Fermi projector. Let $\mathscr{T}_{z}$ and $L_{\eta,j}$ be the operators defined respectively in (3.17) and (3.9). Then, we have that:
@@
(i) & (ii) Note that the operator $J_{i}^{\mathrm{conv}}\,L_{\eta,j}$ is periodic and bounded since both $J_{i}^{\mathrm{conv}}$ and $P,X_{j}]$ are short-range by Lemmas 3.4 and 3.5. Thus, Lemma 3.2(i) ensures that $\sigma_{ij}^{\mathrm{s},\mathrm{conv}}$ is well-defined. On the other hand, for the proper spin conductivity $\sigma_{ij}^{\mathrm{s},\mathrm{prop}}$ we have that $\tau(X_{i}\mathscr{T}_{z}\,L_{\eta,j})=0$ in view of Proposition 3.16(ii), thus only $\mathrm{i}[H,X_{i}]S_{z}L_{\eta,j}\in\mathcal{P}(\mathcal{H})$ contributes to its value. Specifically, we get that
@@
To this end, it is convenient to rewrite the argument of the real part in (3.22) by using that the operator $L_{\eta,j}$ is off-diagonal due to the off-diagonality of $X_{j},P]$. By exploiting the cyclicity of the trace per unit volume, we have that
@@
where we have used repeatedly the cyclicity of the trace per unit volume, and in the last equality we have exploited that $S_{z},[X_{j},P]]=-[P,[S_{z},X_{j}]]-[X_{j},[P,S_{z}]]=[[P,S_{z}],X_{j}]$. By taking the real part of the second term on the last line of (3.24), we obtain the term in (3.19); while in the last line of (3.24) the real part of the first summand gives the third term in the expression for $\Sigma^{\mathrm{snc}}_{ij}$, see (3.20). Next, we collect the remaining terms contributing to the spin conductivity:
@@
(iii) follows immediately from (ii) by using that $\Sigma^{\mathrm{snc}}_{ij}=0$ due to $H,S_{z}]=0=[P,S_{z}]$. ∎
@@
Since Remark 3.10(ii) ensures that the linear response operator $P_{1}$ computed via Kubo’s formula agrees with $\Pi_{1}$ given by the NEASS approach in [MaPaTe], it is clear that a priori these two methods yield the same formulas for the spin conductivies. Here, we have preferred to rewrite the spin-commuting term $\Sigma^{\mathrm{sc}}_{ij}$ in (3.19) by having the operator $S_{z}$ at the first place in the argument of the trace per unit volume, while in [MaPaTe, Theorem 5.6] $S_{z}$ appears nested in a commutator involving $X_{i},P]$ and $X_{j},P]$.